If m, n and p are the roots of the polynomial x3+5x-8 and sum of the roots is 0

0, then find the value of m3+n3+p3.



.



1
Expert's answer
2021-10-04T19:25:14-0400
"m+n+p=-\dfrac{0}{1}=0"

"mn+np+mp=\dfrac{5}{1}=5"

"mnp=-\dfrac{8}{1}=-8"

"(m+n+p)^3=m^3+3m^2(n+p)+3m(n+p)^2+(n+p)^3"

"=m^3+3m^2n+3m^2p+3mn^2+6mnp+3mp^2"

"+n^3+3n^2p+3np^2+p^3"

"=(m^3+n^3+p^3)-3mnp+3m(mn+mp+np)"

"+3n(mn+mp+np)+3p(mn+mp+np)"

"=(m^3+n^3+p^3)-3(-8)+3m(5)+3n(5)"

"+3p(5)=(m^3+n^3+p^3)+24+15(m+n+p)"

"=(m^3+n^3+p^3)+24+15(0)="

"=(m^3+n^3+p^3)+24=(0)^3"

Then


"m^3+n^3+p^3=-24"